OpenStudy (brandonruchin):

Y=2/3x-1/3 X=1/4y+17/4 Solve using substitution The answer is (5,3) I just have no idea how to get that

OpenStudy (mpanderla):

The easiest we to do this question is plug the first equation which is y=... To the second equation for the variable of y. So instead of X=1/4y+17/4, it would be X=1/4(2/3x-1/3)+17/4. Solve for x by rearranging the equation putting all X's on one side and all other numbers on the other side. After gathering like terms. And moving terms. Ultimately solving for X then to solve for the y coordinate plug in the value of X you got back into the original equation that is Y=2/3x-1/3. The answer as you mentioned after doing all of the required steps should be (5,3).